Circular No. 6959

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

    URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1998 M5

F. Shelly reports the discovery of another comet by LINEAR. The following CCD astrometry has been reported:

1998 UT α_{2000} δ_{2000} m_{1} Observer
June 30.30998 23^{h}10^{m}12.^{s}61 +17^{o}49′45′′.4 15.9 Shelly
30.33430 23 10 12.04 +17 50 14.7 16.3  _{”}
30.34652 23 10 11.72 +17 50 28.4 15.9  _{”}
July 1.05806 23 09 51.46 +18 04 34.9 13.8 Korlević
1.05994 23 09 51.41 +18 04 37.4 13.9  _{”}
1.06163 23 09 51.38 +18 04 39.9 13.8  _{”}
1.02361 23 09 52.57 +18 03 54.4  Šarounová
1.02569 23 09 52.50 +18 03 57.0 13.9  _{”}
1.02683 23 09 52.47 +18 03 58.2  _{”}
1.19057 23 09 47.90 +18 07 15.6 14.1 Griffin
1.22039 23 09 46.93 +18 07 49.8 14.2  _{”}
1.24346 23 09 46.24 +18 08 17.6 14.2  _{”}
1.35347 23 09 42.98 +18 10 28.8  Rogers
1.36736 23 09 42.69 +18 10 45.6  _{”}
1.38125 23 09 42.22 +18 11 02.2 14.1  _{”}
1.38194 23 09 42.24 +18 11 03.1 14.7 Offutt
1.39616 23 09 41.73 +18 11 19.8 14.7 Hug
1.40524 23 09 41.45 +18 11 30.5 14.2  _{”}
1.40957 23 09 41.35 +18 11 36.2 14.8 Offutt
1.41681 23 09 41.18 +18 11 45.6 14.8  _{”}

F. Shelly, M. Blythe, and M. Bezpalko (Socorro). 1-m f/2.15 reflector. Measurers J. Stuart, H. Viggh, and R. Sayer. Tail ∼ 40′′ long toward south-southwest.

K. Korlević, V. Jelic, and R. Radonic (Višnjan). 0.4-m f/4.3 reflector. 30′′

coma; tail 120′′ in p.a. 150^{o}. L. Šarounová (Ondřejov). 0.65-m f/3.6 reflector. Coma diameter 30′′; m_{2}

= 15.4; tail at least 4′ long in p.a. 155^{o}. I. P. Griffin (Cocoa). 0.3-m reflector. Coma diameter at least 45′′. J. E. Rogers (Camarillo). 0.25-m Schmidt-Cassegrain reflector. G. Hug (Eskridge). 0.30-m Schmidt-Cassegrain reflector. W. Offutt (Cloudcroft). 0.60-m f/7 Ritchey-Chrétien reflector. Coma di-

ameter 23′′ × 30′′ (wider along northeast-southwest axis); 30′′-wide tail extends 41′′ in p.a. 206^{o}.

1998 July 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green